### v1

# FSFM: A Generalizable Face Security Foundation Model via Self-Supervised Facial Representation Learning

Abstract
This work asks: with abundant, unlabeled real faces, how to learn a robust and transferable facial representation that boosts various face security tasks with respect to generalization performance? We make the first attempt and propose a self-supervised pretraining framework to learn fundamental representations of real face images, FSFM, that leverages the synergy between masked image modeling (MIM) and instance discrimination (ID). We explore various facial masking strategies for MIM and present a simple yet powerful CRFR-P masking, which explicitly forces the model to capture meaningful intra-region Consistency and challenging inter-region Coherency. Furthermore, we devise the ID network that naturally couples with MIM to establish underlying local-to-global Correspondence via tailored self-distillation. These three learning objectives, namely 3C, empower encoding both local features and global semantics of real faces. After pretraining, a vanilla ViT serves as a universal vision Foundation Model for downstream Face Security tasks: cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion facial forgery detection. Extensive experiments on 10 public datasets demonstrate that our model transfers better than supervised pretraining, visual and facial self-supervised learning arts, and even outperforms task-specialized SOTA methods.
1 Introduction
The human face plays a crucial role in daily life and computer vision. In parallel, the facial security landscape suffers from diverse digital and physical manipulations, notably face forgery and presentation attacks. Face forgery edits digital pixels while keeping realistic visual quality. With significant advances in generative models [60, 34, 47], the abuse of this technology, aka deepfakes, could lead to severe trust crises. Presentation attacks use physical media, e.g., printed photos, image/video replays, or 3D masks, to spoof face recognition systems, risking real-life applications like face unlocking and payment. Thus, both academia and industry have been devoted to combating face forgery and presentation attacks with dedicated tasks: Deepfake Detection (DfD), Face Anti-Spoofing (FAS), and recently raised, Diffusion Facial Forgery detection (DiFF).
Current DfD and FAS methods aim to improve generalization to unseen datasets within their respective tasks. DfD methods mainly focus on digital forgery patterns, such as spatiotemporal inconsistency [108, 21, 3], region artifacts [77, 48, 113], and forgery augmentations [113, 77, 3, 89]. FAS methods [50, 134, 62, 133, 92] incorporate domain generalization (DG) techniques to discern physical spoof cues, e.g., paper textures and screen moire patterns. Given these incompatible digital and physical features, most studies treat DfD and FAS as separate face security tasks, i.e., task-specific and hard to transfer. Besides, until now, most DfD and FAS methods still follow fully supervised learning with diverse backbones [22, 95, 28, 43] that are initiated from scratch or supervised ImageNet (natural images) [25] pretraining. However, fully supervised learning requires large-scale annotated data or generative augmentations, which impose expensive resource costs and limit scalability. Moreover, initial weights that lack facial representations may impede capability and generality for face-related tasks [7, 130].
In contrast to supervised learning, self-supervised learning (SSL) takes pretext tasks for pretraining on unlabeled data, where masked image modeling (MIM) [45, 114, 5] and instance discrimination (ID, including contrastive learning [14, 44, 16] and distillation-based [36, 15, 11]) have proven superior performance in various vision tasks. As recent studies [82, 135, 81] suggest that MIM and ID are complementary, SOTA SSL methods [96, 2, 55, 46, 17, 128] combine them to improve the visual representation of natural images. However, SSL progress for facial pretraining, especially security tasks, remains limited, raising Q1: how can face security tasks benefit from self-supervised pretraining to learn scalable and generic representations?
Although existing works introduce SSL to face security tasks, most focus on specific forgery or spoof patterns rather than transferable representations. In DfD, some methods [12, 13, 89, 61] synthesize forgeries from real faces to simulate forgery artifacts like blending, vulnerable to unknown forgery or spoof. Others [31, 41, 127] rely on audio-video datasets and multi-modal information, limiting their scalability. The latest work [123] combines MIM and ID to learn temporal consistency [136] from real face videos, but intra-frame representations fail for image-only forgery and replay videos with consistent temporal features. Conversely, FAS methods employ SSL to mine spoof cues using domain positives [73], domain-invariant semantics [129], or domain alignment [72], yet this FAS domain knowledge contributes less to detecting deepfake and diffusion forgery.
Recent advances in facial pretraining [130, 8, 7, 74, 32] demonstrate that task-agnostic facial representations transfer effectively to several face analysis tasks, such as expression and attribute recognition. However, these methods focus primarily on salient facial features that forgery and spoof faces also exhibit well, i.e., overlook “realness” representations that go beyond normal facial analysis, thus hindering their extrapolation to face security tasks [130]. Furthermore, these methods commonly adopt intra-dataset evaluations for downstream tasks, whereas face security tasks call for cross-dataset generalization. These concerns raise Q2: how to learn universal facial representations from real faces that transfer well to various face security tasks and improve downstream generalization?
To bridge the above gap, instead of task-specific or unified [122, 24, 121] supervised learning, we propose learning the intrinsic properties of unlabeled real face images through SSL. Specifically, we introduce a simple yet effective CRFR-P facial masking strategy into a masked autoencoder [45] that directs attention to pixel-level contexts and region-level correlations, yielding a meaningful and challenging MIM task. By Covering a Random Facial Region (e.g., eyes, nose) and Proportionally masking other regions, our CRFR-P promotes inter-region coherency and intra-region consistency in facial representations. For reliable face-level semantic alignment, we further integrate an ID network with MIM to establish local-to-global correspondence via elaborate self-distillation: the CRFR-P masked online view introduces spatial variance, the unmasked target view retains complete semantics, and Siamese representation decoders form a disentangled space. Thus, our framework, FSFM, leverages the synergy between MIM and ID to empower both local and global perception, learns a universal representation of real faces, transfers well to various face security tasks, and achieves superior generalization in cross-dataset DfD, cross-domain FAS, and unseen DiFF.
Our main contributions are:
We propose a self-supervised pretraining framework to learn a fundamental representation of real faces, FSFM, which fully leverages the advantages of masked image modeling and instance discrimination for both local context perception and global semantic alignment.
We design a simple yet powerful CRFR-P facial masking strategy for MIM to enhance meaningful intra-region Consistency and capture challenging inter-region Coherency, and we develop the ID network coupled with MIM to establish local-to-global Correspondence through elaborate self-distillation—namely, 3C.
We use a vanilla ViT as the encoder to learn transferable and robust facial representations; after pretraining, it serves as a universal vision foundation model for various downstream face security tasks, while previous methods adopt diverse backbones that vary from specific tasks.
We conduct extensive experiments on prevalent face security tasks, including cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion facial forgery detection. Results on 10 public datasets show that our model generalizes better than supervised practices, visual and facial self-supervised pretraining arts, and notably, outperforms task-specialized SOTA methods. Moreover, FSFM benefits from pretraining with more abundantly unlabeled real faces (a free lunch).
2 Related work
Visual representation learning Masked image modeling (MIM) and instance discrimination (ID) are predominant pretext tasks for generative and discriminative self-supervised learning (SSL), respectively. MIM masks parts of an image and takes visible portions to restore the masked content, such as pixels [45, 114], visual tokens [5], or auxiliary features [109]. With a high masking ratio, the tokenizer-free MAE [45] achieves efficient pretraining. Follow-up studies [63, 57, 88, 101, 78] highlight masking strategies and region-based learning. ID employs Siamese networks to pull positive pairs closer, mainly including contrastive learning (CL) and distillation (DIS) paradigms to avoid collapse. CL [14, 44, 16] simultaneously pushes negative pairs away. DIS [36, 15, 11] is negative-free and adopts asymmetric architecture to align representations between two branches. As recent studies [82, 135, 81] suggest that MIM and ID complement each other, cutting-edge SSL methods increasingly combine MIM with ID via Siamese designs [1, 96, 2, 30], including CL [119, 55, 64, 46] and DIS [131, 17, 4, 128], for natural vision representations.
Facial representation learning Several works [93, 97, 79, 100, 66] explore facial SSL to reduce overfitting and improve performance, but they remain tailored for a single task, i.e., task-specific. VLP-based FaRL [130] leverages web image-text CL with MIM to transfer across various facial tasks, but it requires image-text pairs and additional computation overhead, and the text captions tend to describe context rather than face details. Recent SSL efforts aim to learn general facial representations by MIM [8, 103], CL [7, 74, 103], DIS [32, 103] for various downstream tasks, such as expression [8, 7, 74, 32], attribute [8, 32], and AU [7, 74] recognition, alignment [7, 103, 32], etc. Despite excelling in normal face analysis, these methods lack generality for face security tasks as they ignore facial ’realness’ representations.
3 FSFM framework
We now describe the proposed FSFM pretraining framework in Fig. 1, which aims to learn intrinsic and transferable representations of real faces. The pretext tasks for self-supervised learning (SSL) include masked image modeling (MIM) and instance discrimination (ID). Driven by our CRFR-P masking strategy (Sec. 3.1), the MIM network (), a naive masked autoencoder (MAE) [45], reconstructs the masked face to capture intra-region consistency and enforce inter-region coherency (Sec. 3.2). Meanwhile, the ID network employs the MIM encoder in the online branch () and establishes local-to-global correspondence, where the target branch () guides global semantics (Sec. 3.3). These three objectives, termed 3C, endow the online encoder with pixel-level context perceptiveness, region-level relation awareness, and instance-level face invariance.
3.1 CRFR-P masking
The mask sampling strategy plays a key role in MIM for both representation quality and downstream performance, and also provides the local view for our ID network.
Motivation Random masking with a high mask ratio is widely used in natural [45, 114] and facial [130, 103] MIM, but it lacks domain-specific knowledge, limiting pretrained facial models. Given that human faces, our sole focus, comprise well-defined parts with distinct textures, we turn to explicitly segment the facial semantics for reasonable and efficient mask sampling rather than additional learning modules [63, 57, 88, 101]. Similarly, MARLIN [8] uses an off-the-shelf face parser to divide facial parts and introduces the Fasking mask for face video MIM. We adapt Fasking to images as Fasking-I in Fig. 2 (b): it divides facial parts into left-eye, right-eye, nose, mouth, hair, skin, background, prioritizing masking non-skin and non-background regions. However, Fasking-I struggles to capture sufficient details solely from background and skin patches, rendering it unsuitable for face security tasks. Motivated by FACS [29] and psychological studies [85, 42], we explore more effective masking strategies to learn local facial representations.
Intuition of intra-region consistency and inter-region coherency. As shown in Fig. 2, random masking and Fasking-I are prone to fully mask small but informative regions (e.g., eyes), hindering accurate learning of textures therein (e.g., pupils, iris). For intra-region consistency, we introduce FRP (Facial Region Proportional) masking that randomly masks an equal portion of patches in each region. FRP ensures all regions have visible patches, thus enabling attention to the same region when restoring masked patches. However, a potential shortcut, restoring a region’s masked patches directly from its unmasked patches, may lead to a trivial reconstruction task and ignore cross-region correlations. In this regard, we formulate CRFR-R (Covering a Random Facial Region followed by Random) masking for inter-region coherency. As the fully masked region can only be derived from visible patches outside it, CRFR-R encourages learning its relationship w.r.t. other regions. Yet, the subsequent random masking may still obscure other small informative regions, undermining the desired consistency within them. FRP and CRFR-R, our preliminary masking strategies, have individual limitations but complement each other.
CRFR-P Masking Based on the above insights, we propose CRFR-P (Covering a Random Facial Region followed by Proportional) masking strategy, shown in Fig. 1 and Alg. 1. CRFR-P first divides facial parts into predefined regions using an off-the-shelf face parser. Next, it masks all patches within a randomly selected region {skin, background} and obtains the facial region mask . Then, based on the number of masked patches and the overall masking ratio , it randomly masks an equal portion of patches across the remaining regions to generate the image mask . Finally, it returns both and . CRFR-P naturally resolves the core challenge: simultaneously promoting intra-region consistency and inter-region coherency.
Strength Despite its simple design, CRFR-P masking poses a nontrivial and meaningful facial MIM task by effectively directing attention to critical facial regions with appropriate range and diversity. It not only avoids the trivial solution (shortcut) but also captures the intrinsic properties of real faces. In the supplementary material, we reveal key attention differences when applying different facial masking strategies to naive MAE pretraining.
Input: Real face image , Masking ratio
Output: Image mask , Facial region mask
3.2 MIM for facial region perception
In a nutshell, the MIM network () in FSFM is an MAE [45] model guided by our CRFR-P masking strategy. Below, let denote the full set of non-overlapping patches split from the input face image .
Online Encoder only operates on the visible patches and maps into latent features , where means the element-wise product for masking and selects the visible ones. The online encoder first embeds by a linear projection as patch embeddings, and adds corresponding positional embeddings . It then processes the fused embeddings through a series of transformer blocks to obtain: .
Online Decoder reconstructs the input image pixels. It first combines encoded visible tokens with mask tokens , and adds relative positional embeddings to form the full tokens set . The online decoder consists of another sequence of transformer blocks that receives as input, followed by a linear layer to restore the masked patches: .
MIM Objective Following [45], we take normalized pixels as the reconstruction target and minimize the mean squared error (MSE) loss on masked patches between the prediction and the original one :
where is the number of masked patches. Our CRFR-P masking provides an additional mask , covering all patches in a randomly selected facial region . To enhance inter-region coherency and prevent trivial solutions, we add another reconstruction loss only on the masked patches of the facial region :
where is the number of patches in , and . The overall MIM objective updates the MAE network () as a weighted sum:
3.3 ID for local-to-global self-distillation
In a nutshell, the ID network in FSFM employs symmetric designs between the online and target branches w.r.t. the encoder and representation decoder, as well as asymmetric designs w.r.t. the input view, projection head, negative-free loss, and model update rate. These designs focus on more precise and reliable semantic alignment for face security tasks, distinguishing our method from previous works that also integrate ID into MIM or degraded image input.
Target Encoder takes full patches as the target view to produce target latent features that guide the online encoder in learning holistic representations. Among the target view options in Fig. 3, using full patches enables complete facial semantic embedding for local-to-global representation learning. Thus, the target encoder serves as a teacher that shares the same structure as the student . Similarly, with positional embeddings of full patches, yields global embeddings: .
Online and Target Rep Decoder The online rep decoder maps the full tokens to online representations . Unlike the online decoder , which restores pixel values, reconstructs the masked token representations to align with the uncorrupted target, using significantly shallower transformer blocks followed by a linear layer that predicts features. The token features are output by a mean pooling as the online representations: . In the target branch, the momentum encoder is updated by past iterations of the online encoder , which also serves MIM. If is directly matched with target embeddings , the model may struggle to fit high-level target features while restoring low-level pixels. Thus, a target rep decoder with the same structure as is added to represent the target features in the same disentangled space: .
ID Objective Following [36, 15, 16], we use a projector and a predictor to transform into a lower-dimensional vector , and apply only a projector to to obtain . We then minimize the negative cosine similarity [15] between the -normalized and :
where is a stop-gradient, i.e., gradients are only calculated w.r.t. the online branch (). The parameters of the target branch () are updated via an exponential moving average (EMA) [36]: . Note that is asymmetric due to the different input views of the online and target branches, unlike the symmetrized designs [36, 15, 16, 11].
3.4 Overall pretraining objective
4 Experiments
In downstream face security tasks, including deepfake detection (DfD, Sec. 4.2), face anti-spoofing (FAS, Sec. 4.3), and diffusion facial forgery detection (DiFF, Sec. 4.4), we demonstrate the effectiveness of FSFM by asking: Q1: Does our facial representation generalize better than common weight initialization practices? Q2: Is our method superior to other SSL methods for natural vision or normal facial analysis? Q3: Can our pretrained model outperform SOTA specialized methods with only simple fine-tuning? We further ablate FSFM (Sec. 4.5) and present visualizations (Sec. 4.6) to ascertain our contributions. We include more pretraining, finetuning details, ablations, and visualizations in the supplementary material.
| Method | Pretrain or Init | Train Set | Test Set Video-level AUC(%) | Avg. Ours | Method | Pretrain or Init | Train Set | Test Set Frame-level AUC(%) | Avg. Ours | ||||||
| CDFV2 | DFDC | DFDCP | WDF | CDFV2 | DFDC | DFDCP | WDF | ||||||||
| Base model | Base model | ||||||||||||||
| Xception [22] | Sup(IN) | FF++ | 76.39 | 70.62 | 72.24 | 76.11 | 14.0↑ | Xception [22] | Sup(IN) | FF++ | 69.52 | 68.20 | 68.94 | 68.83 | 15.1↑ |
| EfficientNet-B4 [95] | Sup(IN) | FF++ | 79.81 | 71.85 | 66.95 | 76.42 | 14.1↑ | EfficientNet-B4 [95] | Sup(IN) | FF++ | 73.37 | 69.47 | 64.37 | 71.95 | 14.2↑ |
| ViT-B [28] | Scratch | FF++ | 64.08 | 66.73 | 72.62 | 60.36 | 21.9↑ | ViT-B [28] | Scratch | FF++ | 61.14 | 64.27 | 69.00 | 60.68 | 20.2↑ |
| ViT-B [28] | Sup(IN) | FF++ | 86.24 | 74.48 | 82.11 | 81.20 | 6.9↑ | ViT-B [28] | Sup(IN) | FF++ | 77.43 | 71.09 | 74.07 | 75.86 | 9.4↑ |
| MAE [45] ViT-B | SSL(IN) | FF++ | 79.51 | 75.93 | 87.10 | 80.96 | 7.0↑ | MAE [45] ViT-B | SSL(IN) | FF++ | 72.64 | 72.18 | 79.81 | 73.94 | 9.4↑ |
| DINO [11] ViT-B | SSL(IN) | FF++ | 80.47 | 76.90 | 84.64 | 82.06 | 6.9↑ | DINO [11] ViT-B | SSL(IN) | FF++ | 73.88 | 72.78 | 77.31 | 75.08 | 9.2↑ |
| MCF [103] ViT-B | SSL(LFc) | FF++ | 80.25 | 73.61 | 82.55 | 79.79 | 8.8↑ | MCF [103] ViT-B | SSL(LFc) | FF++ | 73.16 | 69.63 | 75.78 | 74.10 | 10.8↑ |
| FSFM ViT-B (Ours) | SSL(VF2) | FF++ | 91.44 | 83.47 | 89.71 | 86.96 | - | FSFM ViT-B (Ours) | SSL(VF2) | FF++ | 85.05 | 80.20 | 85.50 | 85.26 | - |
| SOTA specialized method (Venue) | SOTA specialized method (Venue) | ||||||||||||||
| SBIs [89] (CVPR’22)‡ | Init(IN) | SD | 93.18 | 72.42 | 86.15 | 4.3↑ | OST [13] (NIPS’22)† | Init(IN) | FF++ | 74.80 | 83.30 | 6.2↑ | |||
| RealForensics [41] (CVPR’22) | SSL(LRW) | FF++ | 86.90 | 75.90 | 6.1↑ | FInfer [51] (AAAI’22)† | Init(IN) | FF++ | 70.60 | 70.39 | 69.46 | 15.1↑ | |||
| HCIL [38] (ECCV’22) | Init(IN) | FF++ | 79.00 | 69.21 | 16.5↑ | PEL [37] (AAAI’22)‡ | Init(IN) | FF++ | 69.18 | 63.31 | 67.39 | 16.9↑ | |||
| SeeABLE [61] (ICCV’23)‡ | SSL(SD) | SD | 87.30 | 75.90 | 86.30 | 5.0↑ | SLADD [12] (CVPR’22)† | Init(IN) | SD | 79.70 | 81.80 | 4.5↑ | |||
| TALL [116] (ICCV’23)* | Init(IN) | SD | 90.79 | 76.78 | 3.7↑ | RECCE [9] (CVPR’22)† | Init(IN) | FF++ | 68.71 | 69.06 | 64.31 | 16.1↑ | |||
| AUNet [3] (CVPR’23)† | Init(IN) | SD | 92.77 | 73.82 | 86.16 | 4.0↑ | UIA-ViT [136] (ECCV’22)* | SSL(FF) | FF++ | 82.41 | 75.80 | 6.2↑ | |||
| SLF [21] (CVPR’24) | FTCN | FF++ | 89.00 | 2.4↑ | UAL [112] (MM’23)* | Init(IN) | FF++ | 82.84 | 70.13 | 8.7↑ | |||||
| MLR [48] (CVPR’24)* | FF++ | 91.56 | 75.17 | 73.41 | 7.2↑ | NoiseDF [102] (AAAI’23) | RIDNet | FF++ | 75.89 | 63.89 | 15.3↑ | ||||
| LAA-Net/BI [77] (CVPR’24)‡ | Init(IN) | SD. | 86.28 | 69.69 | 57.13 | 18.3↑ | GS [40] (ICCV’23)‡ | FF++ | 84.97 | 81.65 | 1.8↑ | ||||
| LAA-Net/SBI [77] (CVPR’24)‡ | Init(IN) | SD. | 95.40 | 86.94 | 80.03 | 1.9↑ | UCF [117] (ICCV’23)† | Init(IN) | FF++ | 82.40 | 80.50 | 3.7↑ | |||
| LSDA [118] (CVPR’24)‡ | Init(IN) | FF++ | 91.10 | 77.00 | 3.4↑ | SFDG [104] (CVPR’23)‡ | Init(IN) | FF++ | 75.83 | 73.64 | 69.27 | 10.6↑ | |||
| FPG [113] (MM’24)‡ | Init(IN) | SD. | 94.49 | 74.75 | 87.24 | 2.7↑ | IID [53] (CVPR’23) | CF(WF) | FF++ | 83.80 | 81.23 | 2.8↑ | |||
| NACO [123] (ECCV’24)* | SSL(VC2) | FF++ | 89.50 | 76.70 | 4.4↑ | LSDA [118] (CVPR’24)‡ | Init(IN) | FF++ | 83.00 | 73.60 | 81.50 | 4.2↑ | |||
| Abbreviation: Sup-Supervised SSL-Self-Supervied Init-weight initialization Dataset: IN-ImageNet LRW-LipReading in the Wild/facial VC2-VoxCeleb2/facial CF(WF)-CosFace(WebFace)/facial | |||||||||||||||
| Backbone(or modified): † Xception ‡ EfficientNet-B4 * ViT or transformer-based | |||||||||||||||
4.1 Pretraining setups and baselines
Pretraining Data and Preprocessing We use the VGGFace2 [10] dataset (VF2, 3.3M images) for pretraining. We utilize the DLIB [59] for face detection and cropping with a 30% addition margin, and the FACER [130] toolkit for face parsing instead of alignment. Cropped face images are resized to 224×224, and parsing maps are saved as binary stream files for efficient CRFR-P masking.
Model Architecture Our MIM network is a vanilla MAE [45] with ViT-B/16 as the default encoder . In the ID network: rep decoders, and are 2-layer ViT blocks with the same width as the encoder; the projector and predictor are 2-layer MLPs following BYOL [36]. After pretraining, we only use as the backbone and add a task-specific head for downstream finetuning.
Implementation We set the mask ratio to 0.75. We use no data augmentation (not even crop or flip used in [45]) during pretraining. We empirically set the loss weights , to 0.007, 0.1, respectively. The EMA momentum coefficient follows [36]. We pretrain our model from scratch for 400 epochs. Other setups follow MAE [45] defaults.
Pretrained Base Models To answer Q1 and Q2, we include the following pretrained models based on availability (official weights), fairness (vision-only ViT-B/16 backbone), and relevance: 1) ViT-B Scratch [28], to discern pretraining effects versus the backbone itself; 2) ViT-B Sup(IN) [28], supervised ImageNet pretraining, the most common weight initialization for face security tasks; 3) MAE SSL(IN) [45], our MIM network, the baseline across all experiments, including ablations; 4) DINO SSL(IN) [11], a self-distillation method for ID, which also encourages local-to-global correspondence; 5) MCF SSL(LFc) [103], a SOTA facial representation model for face analysis tasks, which also combines MIM and ID and is pretrained on the large-scale LAION-FACE-cropped dataset (20M images). For downstream tasks, we compare our FSFM with these models by initializing the backbone with their pretrained weights and keeping other settings identical. We employ end-to-end fine-tuning for all models, facilitating a fair comparison with SOTA specialized methods to answer Q3.
4.2 Deepfake detection
| Method | Pretrain or Init | DG FAS Technique | OCI→M | OMI→C | OCM→I | ICM→O | Avg. | |||||||||
| DM | AL | CL | ML | PL | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | ||
| Base model | ||||||||||||||||
| ViT-B [28] | Scratch | 15.37 | 90.73 | 35.37 | 68.23 | 14.75 | 94.18 | 31.65 | 71.55 | 24.28 | 81.17 | |||||
| ViT-B [28](@ECCV’22 [54]) | Sup(IN) | 3.52 | 98.74 | 2.42 | 99.52 | 8.45 | 96.91 | 11.86 | 94.62 | 6.56 | 97.44 | |||||
| MAE [45] ViT-B | SSL(LN) | 10.32 | 94.87 | 15.91 | 89.96 | 15.54 | 91.13 | 16.51 | 90.29 | 14.57 | 91.56 | |||||
| DINO [11] ViT-B | SSL(LN) | 6.73 | 97.15 | 13.44 | 93.90 | 14.27 | 93.56 | 15.55 | 90.99 | 12.50 | 93.90 | |||||
| MCF [103] ViT-B | SSL(LFc) | 4.00 | 98.84 | 8.46 | 96.90 | 8.02 | 97.39 | 10.70 | 95.64 | 7.80 | 97.19 | |||||
| FSFM ViT-B (Ours) | SSL(VF2) | 3.78 | 99.15 | 3.16 | 99.41 | 4.63 | 99.03 | 7.68 | 97.11 | 4.81 | 98.68 | |||||
| SOTA specialized method (Venue) | ||||||||||||||||
| MADDG [87] (CVPR’19)† | Scratch | ✓ | ✓ | ✓ | 17.69 | 88.06 | 24.50 | 84.51 | 22.19 | 84.99 | 27.98 | 80.02 | 23.09 | 84.40 | ||
| SSDG-R [56] (CVPR’20)# | Init(IN) | ✓ | ✓ | 7.38 | 97.17 | 10.44 | 95.94 | 11.71 | 96.59 | 15.61 | 91.54 | 11.29 | 95.31 | |||
| AMEL [132] (MM’22)† | Scratch | ✓ | ✓ | 10.23 | 96.62 | 11.88 | 94.39 | 18.60 | 88.79 | 11.31 | 93.96 | 13.01 | 93.44 | |||
| FGHV [70] (AAAI’22)‡ | Depth Net | ✓ | ✓ | 9.17 | 96.92 | 12.47 | 93.47 | 16.29 | 90.11 | 13.58 | 93.55 | 12.88 | 93.51 | |||
| PatchNet [99] (CVPR’22)# | Init(IN) | ✓ | 7.10 | 98.46 | 11.33 | 94.58 | 13.40 | 95.67 | 11.82 | 95.07 | 10.91 | 95.95 | ||||
| SSAN-R [107] (CVPR’22)# | Init(IN) | ✓ | ✓ | ✓ | 6.67 | 98.75 | 10.00 | 96.67 | 8.88 | 96.79 | 13.72 | 93.63 | 9.82 | 96.46 | ||
| UDG-FAS [73] (ICCV’23)# | SSL(LOO) | ✓ | 7.14 | 97.31 | 11.44 | 95.59 | 6.28 | 98.61 | 12.18 | 94.36 | 9.26 | 96.47 | ||||
| UDG-FAS [73] (ICCV’23)# | Init(IN) | ✓ | 5.95 | 98.47 | 9.82 | 96.76 | 5.86 | 98.62 | 10.97 | 95.36 | 8.15 | 97.30 | ||||
| IADG [133] (CVPR’23)† | Scratch | ✓ | ✓ | 5.41 | 98.19 | 8.70 | 96.44 | 10.62 | 94.50 | 8.86 | 97.14 | 8.40 | 96.57 | |||
| SAFAS [92] (CVPR’23)# | Init(IN) | ✓ | 5.95 | 96.55 | 8.78 | 95.37 | 6.58 | 97.54 | 10.00 | 96.23 | 7.83 | 96.42 | ||||
| GAC-FAS [62] (CVPR’24)# | Init(IN) | 5.00 | 97.56 | 8.20 | 95.16 | 4.29 | 98.87 | 8.60 | 97.16 | 6.52 | 97.19 | |||||
| HPDR [50] (CVPR’24)‡ | Depth Net | ✓ | 4.58 | 96.02 | 11.30 | 94.42 | 11.26 | 92.49 | 9.93 | 95.26 | 9.27 | 94.55 | ||||
| TTDG [134] (CVPR’24)† | Scratch | ✓ | ✓ | 7.91 | 96.83 | 8.14 | 96.49 | 6.50 | 97.98 | 10.00 | 95.70 | 8.14 | 96.75 | |||
| Abbreviation: Sup-Supervised SSL-Self-Supervied Init-weight initialization IN-ImageNet DG-Domain Generalization | ||||||||||||||||
| DG FAS Technique: DM-Depth Maps AL-Adversarial Learning CL-Contrastive Learning (or triplet, similarity loss) ML-Meta Learning PL-Prototype Learning | ||||||||||||||||
| Backbone(or modified): † CNN-based network defined in MADDG ‡ Dense Net # ResNet-18 | ||||||||||||||||
Setting To evaluate the generalizability of our method across diverse DfD scenarios, we follow the challenging cross-dataset setup. Specifically, we train one detector on the FaceForensics++ (FF++, c23/HQ version) [84] dataset and test it on unseen datasets: CelebDF-v2 (CDFv2) [65], Deepfake Detection Challenge (DFDC) [27], Deepfake Detection Challenge preview (DFDCp) [26], and Wild Deepfake (WDF) [137]. For a fair comparison, we report the Area Under Curve (AUC) at both frame-level and video-level, the most widely used metric for DfD, in Tab. 1.
Comparison with Base Models As shown in Tab. 1, our FSFM significantly outperforms all base models on unseen deepfakes at both frame and video levels: 1) It surpasses Sup(IN) baselines, including ViT-B, Xception, and EfficientNet-B4, which are common practices for DfD. This suggests that our pretrained model provides a strong initialization for the detector. 2) MIM-based MAE and ID-based DINO show comparable performance but vary across different datasets, mainly because MIM focuses on local patterns while ID operates globally [135]. FSFM outperforms both MAE and DINO, indicating the effectiveness of learning both local and global facial representations. 3) Notably, FSFM, pretrained on 3M VF2, consistently surpasses MCF, which is pretrained on 20M facial images and also incorporates MIM and ID. Although MCF achieves SOTA performance in face analysis tasks, its generalization to DfD is even worse than ViT, MAE, and DINO pretrained on natural images. 4) These comparisons demonstrate that FSFM effectively learns fundamental real face representations that are sensitive to deepfakes and generalizable for detection.
Comparison with SOTA Specialized Methods As shown by Avg.Ours in Tab. 1, FSFM consistently outperforms all counterparts, achieving the best overall performance. Our model surpasses both transformer-based and SSL-based competitors that are task-specific. In particular, our method outperforms SSL-based NACO, a most recent work [123] that also learns consistent representations of real face videos, and it further combines ViT-B with CNN. Our model also performs best across unseen datasets except in video-level testing on CDFV2, where some models trained on synthetic data show slightly better results. These methods (AUNEt, FPG, and LAA-Net with SBIs) simulate pseudo-fake features, especially blending artifacts similar to face-swapping in CDFV2, but struggle with forgeries lacking such clues. Moreover, our method yields notable gains on more challenging DFDC and WDF datasets, which contain diverse unknown real-world manipulations. Remarkably, our FSFM, a vanilla ViT-B base model, achieves SOTA performance without any specialized network modules or tailored data generation for deepfake detection.
4.3 Face anti-spoofing
Setting To evaluate the transferability of our method for FAS under significant domain shifts, we use four widely used benchmark datasets: MSU-MFSD (M) [110], CASIA-FASD (C) [126], Idiap Replay-Attack (I) [20], and OULU-NPU (O) [6]. We treat each dataset as the target domain and apply the leave-one-out (LOO) cross-domain evaluation. As prior work [54] also finetunes vanilla ViT-B for this protocol, we follow its 0-shot setting by initializing weights from our FSFM and base models. We report mean HTER (Half Total Error Rate) and AUC over 5 runs in Tab. 2.
| Method | Pretrain or Init | Test Subset (AUC%↑) | Avg. w/o FF++ | ||||
| FF++ | T2I | I2I | FS | FE | |||
| ViT-B [28] | Scratch | 92.02 | 62.19 | 69.99 | 60.87 | 67.30 | 65.09 |
| ViT-B [28] | Sup(IN) | 99.15 | 33.38 | 35.83 | 52.20 | 55.42 | 44.21 |
| MAE [45] ViT-B | SSL(IN) | 99.25 | 33.01 | 32.88 | 47.77 | 58.70 | 43.09 |
| DINO [11] ViT-B | SSL(IN) | 99.30 | 33.85 | 36.02 | 60.37 | 63.18 | 48.35 |
| MCF [103] ViT-B | SSL(LFc) | 99.39 | 39.09 | 38.67 | 34.35 | 56.02 | 42.03 |
| FSFM ViT-B | SSL(FF++_o) | 99.31 | 61.74 | 71.91 | 71.31 | 78.98 | 70.99 |
Comparison with Base Models Our FSFM achieves the best average performance among base models, with several key observations: 1) Simple finetuning on ViT-B Sup(IN) enhances FAS generalization, as noted in [54, 134]. 2) General SSL methods, including MIM-based MAE [45] and ID-based DINO [11], perform worse than ViT-B Sup(IN), with MAE showing notable declines, echoing the claims in [120]. 3) With large-scale facial data for SSL, MCF outperforms MAE and DINO for FAS. However, similar to DfD, it still falls short of the ViT-B Sup(IN) baseline in average metrics, once again underscoring the gap between normal face analysis and security tasks. 4) Overall, our FSFM improves the generalizability of ViT for cross-domain FAS, effectively modeling credible features of live (real) faces.
Comparison with SOTA Specialized Methods Our FSFM ViT-B exhibits superior performance over domain generalization (DG) FAS methods using various CNNs. Notably, this comparison adheres to the baseline [54]: although supplementary data (CelebA-Spoof [124]) is included for fine-tuning ViT-B, no auxiliary supervision or DG techniques, e.g., depth map (DM) or adversarial learning (AL), are employed—only a standard cross-entropy loss. Additionally, we follow other ViT-based FAS methods [134, 67, 106] without including CelebA-Spoof, with comparisons provided in the supplementary material.
4.4 Diffusion facial forgery detection
Setting To further investigate the adaptability of our method against emerging unknown facial forgeries, we adopt cross-distribution testing using the recently released DiFF [19] benchmark. This dataset contains high-quality face images synthesized by 13 SOTA diffusion models across four subsets: T2I (Text-to-Image), I2I (Image-to-Image), FS (Face Swapping), and FE (Face Editing). We train only on the FF++_DeepFake (c23) subset and report AUC results on the DiFF testing subsets. This evaluation is more challenging than typical DfD (Sec. 4.2), as both the unseen manipulations and generative models are significantly different.
Comparison In Tab. 3, our model significantly outperforms other ViT detectors. Pretrained base models perform even worse than the model trained from scratch. This arises from severe overfitting to the deepfake distribution, which hinders discerning diffusion-generated faces. In contrast, our model benefits from fundamental representations of real faces that are not specific to forgery types. The substantial improvement in this new task highlights the robustness of our method in out-of-distribution scenarios.
4.5 Ablation studies
| Component | Deepfake Detection | Face Anti-spoofing | |||||
| F-AUC↑ | V-AUC↑ | HTER↓ | AUC↑ | ||||
| MAE &Masking Strategy (w/o ID Network) | |||||||
| &Random (base) | 74.19 | 79.51 | 19.05 | 87.42 | |||
| &Fasking-I [8] | 73.80 | 78.33 | 17.81 | 87.75 | |||
| &FRP | 75.43 | 81.21 | 17.96 | 87.61 | |||
| &CRFR-R | 75.01 | 80.70 | 18.28 | 87.34 | |||
| &CRFR-P | 76.11 | 81.58 | 17.85 | 88.11 | |||
| ID &Target View (w/ MAE&CRFR-P) | |||||||
| &Visible | 75.54 | 81.50 | 18.22 | 87.95 | |||
| &Masked | 76.35 | 81.86 | 18.41 | 87.77 | |||
| &Full (FSFM) | 76.39 | 82.31 | 17.44 | 88.26 | |||
| Design | Setting | ||||||
| Online&Target Rep Decoder (&) Blocks | 0 & 0 | 75.63 | 81.48 | 18.37 | 86.77 | ||
| 2 & 0 | 75.74 | 81.14 | 18.54 | 87.22 | |||
| 1 & 1 | 75.06 | 80.68 | 18.86 | 87.64 | |||
| 2 & 2 | 76.39 | 82.31 | 17.44 | 88.26 | |||
| 3 & 3 | 75.08 | 80.71 | 17.93 | 87.80 | |||
| Online&Target Data (&) Augmentation | crop+flip&none | 75.93 | 81.54 | 18.24 | 87.04 | ||
| none&crop+flip | 73.39 | 78.80 | 19.13 | 86.11 | |||
| none&none | 76.39 | 82.31 | 17.44 | 88.26 | |||
| Loss for ID | InfoNCE | 75.10 | 80.60 | 18.24 | 87.37 | ||
| MSE ([36]-like) | 74.19 | 79.34 | 18.09 | 88.21 | |||
| Asym. Eq. 4 | 76.39 | 82.31 | 17.44 | 88.26 | |||
| pretraining dara | Size | ||||||
| Data size (images) | FF++_o (0.1M) | 76.39 | 82.31 | 17.44 | 88.26 | ||
| YTF(0.6M) [111] | 79.51 | 83.86 | 16.23 | 93.13 | |||
| VF2 (3.3M) | 84.00 | 87.89 | 4.81 | 98.68 | |||
We conduct extensive ablations to evaluate the effectiveness of each component. In this subsection, unless otherwise stated, we pretrain the ViT-B/16 network on the FF++_o dataset, which includes 0.11M real face images extracted from 720 training and 140 validation videos in the FF++ (c23) Youtube subset [84]. We present the average generalization performance on DfD and FAS in Tab. 4.
Effect of 3C We first evaluate facial masking strategies in the vanilla MAE. Both our preliminary FRP and CRFR-R outperform random masking, validating the significance of intra-region consistency () and inter-region coherency (), respectively. Further, our CRFR-P emerges as the most effective, showing that both and are critical and complementary for robust facial representation. Building on MAE&CRFR-P, we incorporate the ID network with varied target views (Fig. 3). The improved performance with &Full confirms the benefit of local-to-global correspondence () guided by complete facial semantics.
Effect of Key Designs 1) Blocks of & A lightweight 2-layer Rep decoder in both the online and target branches (2&2) achieves a better balance between complexity and generalization, outperforming 1&1 and 3&3. Omitting the Rep decoder (0&0) or adding it only to the target branch (2&0) degrades performance, supporting the importance of bridging feature distribution in a disentangled space. 2) Data Augmentation FSFM performs best without data augmentation (none&none). Unlike many SSL methods, applying simple augmentation in the MIM network (crop+flip&none) or target view (none&crop+flip) reduces performance, indicating that CRFR-P masking introduces sufficient spatial regularization and that preserving target representations from original faces aids robustness. 3) Loss for ID Network Pretraining with asymmetric negative cosine similarity (NCS) i.e. Eq. 4 consistently outperforms the widely used InfoNCE, as the latter pushes different real faces (negative samples) apart that may not be suitable for face security tasks. Besides, our SimSiam-like [15] NCS proves more effective than BYOL-like [36] MSE.
Effect of Data Scaling As expected, pretraining FSFM on a larger dataset yields substantial gains, which suggests that incorporating more diverse real faces in open-ended environments can further improve generalization.
4.6 Qualitative analysis
To illustrate the superiority of the facial features learned by FSFM, we visualize GradCAM [86] activation maps for detecting forgery and spoofing, compared to Sup(IN) and MAE baselines. 1) DfD In Fig. 4 (a), FSFM heatmaps more accurately identify forgery-relevant artifacts in FF++ [84], e.g., mouth-modified F2F and NT. 2) FAS In Fig. 4 (b), FSFM predominantly captures spoof-specific clues in the MCIO cross-domain testings: high-frequency paper textures (M-Paper), photo cut edges (C-Photo), inconsistent reflections between facial regions (I-Print), and moiré patterns in video (O-Replay). These visualizations show that FSFM effectively responds to anomalies diverging from the suggested 3C in real/live faces, shedding light on improving the generalization of face security tasks.
5 Conclusion
In this work, we propose a self-supervised pretraining framework to learn fundamental and general facial representations of real faces, FSFM. We leverage the synergy between masked image modeling and instance discrimination to empower both local and global perception, promoting intra-region consistency, inter-region coherency, and local-to-global correspondence. We show that our FSFM transfers better than previous pretrained models on several face security tasks, including cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion face forgery detection. Notably, as a vanilla ViT, our model surpasses SOTA task-specific methods on generalization performance. Finally, we hope that our work can contribute to the face security-related research.
Supplementary Material
A Overview
This supplementary material provides additional insights, details, and results supporting our FSFM framework, structured as follows:
Facial Masking Strategies in MIM (Sec. B): We delve into the impact of different facial masking strategies on naive MAE, including quantitative and qualitative analyses of attention differences.
Instance Discrimination in FSFM (Sec. C): We highlight the connections and distinctions between our method and prior works that also integrate ID (or Siamese encoder architecture) into MIM (or degraded input).
Implementation Details (Sec. D): Detailed descriptions of hyperparameters, pretraining, and finetuning settings.
Additional Results and Comparisons (Sec. E): Extended experiments comparing FSFM against other models, including ViT-based FAS and the base VLP model, CLIP.
Ablations and Visualizations (Sec. F): Supplementary evidence validating FSFM’s ability to learn robust and transferable facial representations.
B Revealing the secrets of facial masking strategies in MIM
In the main paper, we explore various facial masking strategies for masked image modeling (MIM) in FSFM, with additional visualizations provided in Fig. 5, and validate the effectiveness of CRFR-P masking through ablation studies on downstream face security tasks. However, a critical question remains: how do different facial masking strategies affect the MIM pretrained model or its learned representations of real faces?
To address this, we quantitatively and qualitatively analyze the properties of attention maps. Given that most MIM pretrained models, including ours, are built on the Vision Transformer (ViT) architecture [28], whose main component, the attention mechanism [98], is naturally interpretable [115]. Here, we adopt the naive MAE (FSFM without the ID network) and follow its settings [45] with ViT-B/16 as the encoder and a 75% masking ratio. We conduct self-supervised pretraining on real face images from FF++_o [84] (the default dataset for our ablations). We alter only the masking strategy, encompassing simple random masking, Fasking-I, FRP, CRFR-R, and CRFR-P, and we examine the following aspects of attention heads in the pretrained models: 1) mean attention distance to measure the flow of local and global facial information; 2) KL divergence to investigate the diversity of attention; 3) visualized attention maps to uncover key regions of focus.
B.1 Local or global?
To explore whether the pretrained model attends to faces locally or globally, we calculate the mean attention distance [28] in each attention head across all blocks/layers, as shown in Fig. 6 (top). The model (MAE ViT-B/16 encoder) pretrained with simple random masking tends to focus on local information in the lower blocks and shifts toward global attention in the deeper blocks, similar to the supervised model [28]. Fasking-I primarily aggregates global information as the visible patches predominantly consist of broad backgrounds and skin. FRP also causes large mean attention distances, but they are slightly smaller than those of Fasking-I, mainly because visible patches in FRP are more evenly distributed across all facial regions. CRFR-R fully masks a facial region before applying random masking, which encourages attention to different regions. Consequently, CRFR-R has more global attention in the middle (3rd to 8th) blocks than the simple random masking counterparts. Compared with CRFR-R, CRFR-P masks the remaining regions proportionally instead of randomly, making the 1st block more global w.r.t. the more unmasked regions. Compared with FRP, CRFR-P fully masks a region before applying proportional masking, exposing more patches within other regions at the same masking ratio, thus leading to more local attention than FRP.
Overall, the model pretrained with CRFR-P exhibits well-distributed mean attention distances across all blocks, indicating a synergistic effect of FRP and CRFR-R, enabling appropriate attention to both local and global information.
B.2 Similar or different?
To assess whether the pretrained model focuses on similar or different tokens, we compute the Kullback-Leibler (KL) divergence between attention maps of each head across all blocks, following [115], as shown in Fig. 6 (bottom). As the visible patches are mostly background and skin, the model pretrained with Fasking-I aggregates similar tokens, leading to low KL divergence across all attention heads. Interestingly, we find that proportional masking reduces diversity among attention heads, likely due to its homogeneous presentation of visible tokens, i.e., derived from each facial region. In contrast, covering a random facial region increases attention diversity, as evidenced by higher KL divergence in CRFR-R versus the simple random masking counterparts and CRFR-P versus the FRP counterparts. This suggests that the model is compelled to look at different regions after fully masking a facial region.
Overall, the model pretrained with FRP lacks diversity across attention heads, while CRFR-R shows excessive diversity. Similar to the phenomenon observed in mean attention distance, CRFR-P strikes a balance in KL divergence across different heads, which also seems to act as a co-effect of FRP and CRFR-R counterparts, implicating appropriate attention to different key tokens.
B.3 Key regions?
To uncover which regions of real faces are critical to the pretrained model, we visualize the mean attention map of the last block and overlay it onto the input face in Fig. 7, as the pretext decoder or downstream head follows the last block. We can observe significant attention differences in the salient regions across different masking strategies. At first glance, the attention regions from simple random masking appear to cover the entire face. However, it predominantly highlights the skin, which can be easily recovered from visible neighboring patches, ignoring more challenging key regions. This suggests that the model solves face reconstruction through shortcuts instead of learning meaningful features. Similarly, the attention in Fasking-I is distributed across skin and background regions, as expected. While FRP and CRFR-R activate more attention areas, they still struggle to focus on key facial regions. In contrast, CRFR-P highlights attention across key regions like the eyes, eyebrows, and nose, indicating that the pretrained model tackles the challenge head-on: focusing on meaningful region-level features beyond just low-level pixels of real faces.
In summary, the CRFR-P masking strategy effectively directs attention to critical facial regions with appropriate range and diversity for both intra-region consistency and inter-region coherency, enabling the pretrained facial model to avoid trivial solutions (shortcuts) and capture the intrinsic properties of real faces. Furthermore, we hope this section provides new insights into fundamental face representation.
| Method | Pretrain or Init | DG FAS Technique | OCI→M | OMI→C | OCM→I | ICM→O | Avg. | |||||||||
| DM | AL | CL | ML | PL | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | ||
| Base model | ||||||||||||||||
| ViT-B [28]* | Scratch | 15.25 | 90.07 | 36.98 | 63.94 | 10.75 | 96.09 | 37.50 | 66.03 | 25.12 | 79.03 | |||||
| ViT-B [28]* | Sup.(IN) | 5.00 | 97.29 | 11.51 | 95.61 | 10.60 | 94.03 | 17.35 | 88.35 | 11.12 | 93.82 | |||||
| MAE [45] ViT-B* | SSL.(LN) | 8.17 | 97.37 | 27.91 | 77.90 | 18.02 | 91.54 | 25.36 | 80.76 | 19.86 | 86.89 | |||||
| DINO [11] ViT-B* | SSL.(LN) | 7.92 | 97.09 | 21.28 | 87.79 | 21.35 | 84.48 | 17.44 | 89.75 | 17.00 | 89.78 | |||||
| MCF [103] ViT-B* | SSL.(LFc) | 6.33 | 98.29 | 21.40 | 86.72 | 13.13 | 95.15 | 16.76 | 89.59 | 14.41 | 92.44 | |||||
| FSFM ViT-B (Ours)* | SSL.(VF2) | 6.58 | 97.43 | 4.30 | 99.10 | 14.63 | 90.96 | 10.02 | 96.23 | 8.88 | 95.93 | |||||
| ViT-based specialized method (Venue) | ||||||||||||||||
| ViTranZFAS [33] (IJCB’21)* | Init.(IN) | 10.95 | 95.05 | 14.33 | 92.10 | 16.64 | 85.07 | 15.67 | 89.59 | 14.40 | 90.45 | |||||
| TransFAS [105] (TBIOM’22)* | Init.(IN) | ✓ | 7.08 | 96.69 | 9.81 | 96.13 | 10.12 | 95.53 | 15.52 | 91.10 | 10.63 | 94.86 | ||||
| TTN-S [106] (TIFS’22)* | Init.(IN) | ✓ | 9.58 | 95.79 | 9.81 | 95.07 | 14.15 | 94.06 | 12.64 | 94.20 | 11.55 | 94.78 | ||||
| DiVT-V [67] (WACV’23)* | Init.(IN) | ✓ | ✓ | 10.00 | 96.64 | 14.67 | 93.08 | 5.71 | 97.73 | 18.06 | 90.21 | 12.11 | 94.42 | |||
| TTDG-V [134] (CVPR’24)* | Init.(IN) | ✓ | 4.16 | 98.48 | 7.59 | 98.18 | 9.62 | 98.18 | 10.00 | 96.15 | 7.84 | 97.75 | ||||
| Abbreviation: Sup-Supervised SSL-Self-Supervied Init-weight initialization IN-ImageNet DG-Domain Generalization | ||||||||||||||||
| DG FAS Technique: DM-Depth Maps AL-Adversarial Learning CL-Contrastive Learning (or triplet, similarity loss) ML-Meta Learning PL-Prototype Learning | ||||||||||||||||
| Method | Pretrain or Init | Train Set | Test Set Video-level AUC(%) | Avg. Ours | Method | Pretrain or Init | Train Set | Test Set Frame-level AUC(%) | Avg. Ours | ||||||
| CDFV2 | DFDC | DFDCP | WDF | CDFV2 | DFDC | DFDCP | WDF | ||||||||
| Base model | Base model | ||||||||||||||
| Xception [22] | Sup(IN) | FF++ | 76.39 | 70.62 | 72.24 | 76.11 | 14.0↑ | Xception [22] | Sup(IN) | FF++ | 69.52 | 68.20 | 68.94 | 68.83 | 15.1↑ |
| EfficientNet-B4 [95] | Sup(IN) | FF++ | 79.81 | 71.85 | 66.95 | 76.42 | 14.1↑ | EfficientNet-B4 [95] | Sup(IN) | FF++ | 73.37 | 69.47 | 64.37 | 71.95 | 14.2↑ |
| ViT-B [28] | Scratch | FF++ | 64.08 | 66.73 | 72.62 | 60.36 | 21.9↑ | ViT-B [28] | Scratch | FF++ | 61.14 | 64.27 | 69.00 | 60.68 | 20.2↑ |
| ViT-B [28] | Sup(IN) | FF++ | 86.24 | 74.48 | 82.11 | 81.20 | 6.9↑ | ViT-B [28] | Sup(IN) | FF++ | 77.43 | 71.09 | 74.07 | 75.86 | 9.4↑ |
| MAE [45] ViT-B | SSL(IN) | FF++ | 79.51 | 75.93 | 87.10 | 80.96 | 7.0↑ | MAE [45] ViT-B | SSL(IN) | FF++ | 72.64 | 72.18 | 79.81 | 73.94 | 9.4↑ |
| DINO [11] ViT-B | SSL(IN) | FF++ | 80.47 | 76.90 | 84.64 | 82.06 | 6.9↑ | DINO [11] ViT-B | SSL(IN) | FF++ | 73.88 | 72.78 | 77.31 | 75.08 | 9.2↑ |
| MCF [103] ViT-B | SSL(LFc) | FF++ | 80.25 | 73.61 | 82.55 | 79.79 | 8.8↑ | MCF [103] ViT-B | SSL(LFc) | FF++ | 73.16 | 69.63 | 75.78 | 74.10 | 10.8↑ |
| CLIP [83] ViT-B | VLP(WIT) | FF++ | 78.95 | 73.83 | 82.38 | 78.60 | 9.5↑ | CLIP [83] ViT-B | VLP(WIT) | FF++ | 73.02 | 70.66 | 77.46 | 72.04 | 10.7↑ |
| FSFM ViT-B (Ours) | SSL(VF2) | FF++ | 91.44 | 83.47 | 89.71 | 86.96 | - | FSFM ViT-B (Ours) | SSL(VF2) | FF++ | 85.05 | 80.20 | 85.50 | 85.26 | - |
| Abbreviation: Sup-Supervised SSL-Self-Supervied VLP-Vision Language pretraining Init-weight initialization | |||||||||||||||
| Dataset: IN/1M natural images [25] LFc/20M facial images [103] WIT/400M (natural image, text) pairs [83] VF2/3M facial images [10] | |||||||||||||||
C Connection and analysis of ID in FSFM
We illustrate the relation and distinction between our method and previous works that integrate ID (or Siamese encoder architecture) into MIM (or degraded input). While these hybrid approaches have demonstrated effectiveness in natural vision and face analysis, our empirical studies reveal that face security tasks necessitate more precise and reliable semantic alignment from the ID network. In response, we distinguish key designs such as target view&network structure, data augmentation, and loss function, which support local-to-global correspondences in FSFM.
C.1 Target view and network design
From the perspective of the input view, the online/student branch typically processes visible patches from the masked image, while the target/teacher branch varies across methods. Thus, we incorporate different target views and design paradigms into FSFM, as shown in Fig. 8. (a) Visible patches from a different mask [64, 103, 4, 128]: the online and target encoders produce latent features and for subsequent contrast learning. (b) Masked patches from the same mask [17]: to align the representation of masked patches with the encoded target , a masked representation (rep) decoder predicts from the visible tokens output by the online encoder. This decoder computes cross-attention between masked tokens (as Q) and full tokens (as K and V), following the latent regressor in CAE [17] and resembling the prompting decoder in [49]. (c) Full patches without masking [131, 119, 1, 55, 96]: some methods [131, 119, 1] are decoder-free designs that match visible online features with full target features to compute . Unlike CMAE [55], which introduces a feature/rep decoder after the online encoder, i.e., , we add an additional target rep decoder to compute in a disentangled representation space. i.e., Siamese rep decoders. This design further reduces the gap in distribution between low-level pixel features and high-level semantic representations.
Based on our ablations (in the main paper) of downstream face security tasks, FSFM performs better when using full patches as the target view alongside Siamese rep decoders. By predicting the entire face representation from visible parts, the ID network aligns global and local views of the same face. In light of this, FSFM structures the encoded space with semantically complete facial representations through ”local-to-global” correspondences [11], which endows the encoder with strong facial discriminability.
C.2 Data augmentation
Most ID methods [14, 44, 16, 36, 15, 11] rely on strong data augmentations, including spatial and color transforms, to prevent trivial solutions. For MIM, applying strong augmentations, such as color enhancements, is suboptimal [45], as masking corruption itself effectively regularizes the pretext task. This is further evident in methods [119, 64, 55] that combine MIM and ID, where only simple augmentations—random size cropping or flipping—are applied to the masked input of the online branch, while strong or simple augmentations are used for the full (unmasked) target view.
In contrast, our FSFM behaves well without any data augmentation in either the online or target branches. This may stem from the semantic integrity preserved in unaugmented inputs, which benefits the learning of global face identity [103], especially in face security tasks where forgery and spoofing cues may be implicit anywhere. Additionally, the proposed CRFR-P masking strategy inherently introduces spatial variance tailored to facial structures, rendering simple augmentations (crop and flip) unnecessary. Consequently, FSFM processes only a single view per face image.
| Method | Pretrain or Init | OCI→M | OMI→C | OCM→I | ICM→O | Avg. | |||||
| HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | HTER↓ | AUC↑ | ||
| Base model | |||||||||||
| ViT-B [28] | Scratch | 15.37 | 90.73 | 35.37 | 68.23 | 14.75 | 94.18 | 31.65 | 71.55 | 24.28 | 81.17 |
| ViT-B [28, 54] | Sup(IN) | 3.52 | 98.74 | 2.42 | 99.52 | 8.45 | 96.91 | 11.86 | 94.62 | 6.56 | 97.44 |
| MAE [45] ViT-B | SSL(LN) | 10.32 | 94.87 | 15.91 | 89.96 | 15.54 | 91.13 | 16.51 | 90.29 | 14.57 | 91.56 |
| DINO [11] ViT-B | SSL(LN) | 6.73 | 97.15 | 13.44 | 93.90 | 14.27 | 93.56 | 15.55 | 90.99 | 12.50 | 93.90 |
| MCF [103] ViT-B | SSL(LFc) | 4.00 | 98.84 | 8.46 | 96.90 | 8.02 | 97.39 | 10.70 | 95.64 | 7.80 | 97.19 |
| CLIP [83] ViT-B | VLP(WIT) | 4.29 | 98.76 | 5.00 | 98.89 | 7.14 | 97.92 | 6.09 | 98.12 | 5.63 | 98.42 |
| FSFM ViT-B (Ours) | SSL(VF2) | 3.78 | 99.15 | 3.16 | 99.41 | 4.63 | 99.03 | 7.68 | 97.11 | 4.81 | 98.68 |
| Abbreviation: Sup-Supervised SSL-Self-Supervied VLP-Vision Language pretraining Init-weight initialization | |||||||||||
| Dataset: IN-ImageNet1K [25] LFc-LAION FACE cropped [103] WIT-WebImageText [83] VF2-VGGFace2 [10] | |||||||||||
C.3 Loss function for ID
We consider two main types of loss functions for ID: contrastive [14, 16] and non-contrastive [36, 15]. Contrastive loss pulls positive views from the same sample together and pushes negative views from different samples apart. We use the widely adopted InfoNCE [80] as the contrastive loss. Non-contrastive loss maximizes the similarity of positive representations only. We use mean squared error (MSE) in BYOL [36] and negative cosine similarity (NCS) in SimSiam [15] as non-contrastive loss, respectively, but in an asymmetric formulation, as detailed in the main paper.
In FSFM, we observe that NCS outperforms the contrastive loss InfoNCE, even though methods combining MIM and ID [64, 119, 55, 103] favor the latter. We speculate this is because, in large-scale pretraining on real faces, the inter-image contrast introduced by negative sample pairs—pushing one real face away from others—does not help our model learn representations that are beneficial for face security tasks. Thus, we adopt asymmetric NCS by default to learn intra-face correspondences by matching the online anchor view with the target view of the same sample.
D More implementation details
D.1 pretraining settings
We set the mask ratio to 0.75, similar to the baseline [45], as our ablation shows that this high ratio is also favorable for our FSFM. We do not use any data augmentation (not even crop and flip used in [45]) and only normalize the input faces during pretraining. We empirically set the loss weights and to 0.007 and 0.1, respectively. The projection and prediction heads are 2-layer MLPs following BYOL [36], with batch normalization (BN) replaced by layer normalization (LN) for our ViT-based architecture. The EMA momentum coefficient for updating the target branch starts from 0.996 and increases with a cosine scheduler, following BYOL [36]. We pretrain our model from scratch for 400 epochs on 4 NVIDIA RTX A6000 GPUs. Other settings follow the defaults in MAE [45]: we use the AdamW [75] optimizer with momentum , ; we apply the linear lr scaling rule [35] with a base learning rate of 1.5e-4; we adopt a cosine decay [76] learning rate schedule with a warmup epoch of 40; we maintain the effective batch size as 4096 = 256 (batch size per GPU) 4 (GPUs) 4 (accumulated gradient iterations).
D.2 Finetuning settings in downstream tasks
For finetuning the ViTs from FSFM and other pretrained models, we adopt identical settings except for weight initialization, detailed below.
Deepfake Detection We use the c23 (HQ) version of FF++ [84] with official train/val splits for finetuning. We sample 128 frames per real video (the original YouTube subset) and 32 frames per forgery video (including Deepfakes, Face2Face, FaceSwap, and NeuralTextures subsets). We follow the official test split in other unseen datasets for testing, including CDFV2 [65], DFDC [27], DFDCp [26], and WDF [137]. We sample 32 frames per testing video. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224×224. As WDF already provides 224×224 facial images, we directly use its test set. We add only one linear layer as the binary classifier after averaging all non-CLS token features. We set the batch size to 64, the base learning rate to 2.5e-4, and the finetuning epochs to 10 (50 for ViT-B Scratch). Other settings adhere to the MAE ImageNet finetuning recipe [45].
Face Anti-Spoofing In the main paper, we adopt the 0-shot MCIO setting (Protocol 1) in [54] and include CelebA-Spoof [124] as supplementary data for FAS finetuning. We set the batch size to 12 for each training domain. We append the MLP head after averaging all non-CLS token features instead of using CLS ones [54], to align with other face security tasks. Additionally, for a fair comparison with other visual-only ViT-based methods, we additionally follow [134] and report the best performance without including the supplementary data, as presented in Sec. E.1.
Diffusion Face Forgery Detection For the training set, we sample 32 frames from each real video (the original YouTube subset) and each forgery video (the Deepfakes subset) from FF++ (c23) [84]. For validation and testing sets, we follow the splits provided by the DiFF benchmark [18]. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224×224. We add one linear layer as the binary classifier after averaging all non-CLS token features. We set the batch size to 256, the base learning rate to 5e-4, and the finetuning epochs to 50. Other settings adhere to the MAE ImageNet finetuning recipe [45].
E Additional experimental results
E.1 Comparison with ViT-based FAS
In a fair comparison with visual-only ViT-based FAS methods, our method also significantly outperforms all base models, as shown in Tab. 5. FSFM surpasses most counterparts and ranks second in average metrics. TTDG-V [134], which introduces test-time domain generalization and explicit optimization goals for FAS, performs better than ours in two out of four target domains (M and I). While optimizing for a specific downstream task is beyond the scope of this study, incorporating special auxiliary supervision or domain generalization (DG) techniques into our pretrained model may further improve its generalization ability for face presentation attack detection.
E.2 Comparison with CLIP
| Method | Pretrain or Init | Test Subset (AUC%↑) | Avg. w/o FF++ | ||||
| FF++ | T2I | I2I | FS | FE | |||
| ViT-B [28] | Scratch | 92.02 | 62.19 | 69.99 | 60.87 | 67.30 | 65.09 |
| ViT-B [28] | Sup(IN) | 99.15 | 33.38 | 35.83 | 52.20 | 55.42 | 44.21 |
| MAE [45] ViT-B | SSL(IN) | 99.25 | 33.01 | 32.88 | 47.77 | 58.70 | 43.09 |
| DINO [11] ViT-B | SSL(IN) | 99.30 | 33.85 | 36.02 | 60.37 | 63.18 | 48.35 |
| MCF [103] ViT-B | SSL(LFc) | 99.39 | 39.09 | 38.67 | 34.35 | 56.02 | 42.03 |
| CLIP [83] ViT-B | VLP(WIT) | 99.33 | 69.63 | 66.25 | 65.23 | 57.07 | 64.54 |
| FSFM ViT-B | SSL(FF++_o) | 99.31 | 61.74 | 71.91 | 71.31 | 78.98 | 70.99 |
Another line of representation learning, vision-language pretraining (VLP), particularly contrastive language-image pretraining (CLIP) [83], has shown remarkable zero-shot and generalization capabilities across diverse downstream tasks. Recent studies have successfully tailored CLIP to specific face security tasks, including deepfake detection [58, 94, 90], face anti-spoofing [91, 69, 52, 39, 71], and diffusion forgery detection [125, 23, 68]. These text-aided methods differ fundamentally from our FSFM, which is vision-only, self-supervised, and task-agnostic. Moreover, VLP demands extensive (image, text) data pairs along with significant computing resources for the additional text encoder. Despite these, we include CLIP as a base vision-language model (VLM) for comparison.
Specifically, we borrow the CLIP image encoder, also a ViT-B/16 but pretrained on 400M (image, text) pairs, and finetune it on downstream face security tasks under the same settings as other base models. We supplement its results on deepfake detection, face anti-spoofing, and diffusion face forgery detection in Tab. 6, Tab. 7, and Tab. 8, respectively. We can observe that CLIP ViT-B transfers better than other base vision models on FAS and DiFF tasks, benefiting from the extensive scale of multi-modal supervision. However, directly applying CLIP ViT-B to DfD exhibits inferior generalization. In contrast, our proposed FSFM also consistently outperforms CLIP ViT-B across downstream face security tasks.
F More ablations and visualizations
F.1 Ablation studies
| Component | Setting | Deepfake Detection | Face Anti-spoofing | ||
| F-AUC↑ | V-AUC↑ | HTER↓ | AUC↑ | ||
| Masking ratio | 0.35 | 73.92 | 79.56 | 15.84 | 90.02 |
| 0.50 | 74.31 | 79.48 | 21.57 | 84.09 | |
| 0.65 | 74.92 | 80.13 | 17.76 | 87.37 | |
| 0.75 | 76.39 | 82.31 | 17.44 | 88.26 | |
| 0.85 | 75.40 | 80.83 | 19.19 | 86.13 | |
| pretraining model | Size | ||||
| Model size (parameters) | ViT-S/16(22M) | 74.80 | 80.20 | 19.32 | 89.13 |
| ViT-B/16(86M) | 76.39 | 82.31 | 17.44 | 88.26 | |
| ViT-L/16(303M) | 77.43 | 83.15 | 16.23 | 93.13 | |
This subsection presents additional ablations. Unless otherwise stated, the default settings follow the main paper.
Effect of Masking Ratio We also examine the impact of different masking ratios for CRFR-P masking on our pretraining framework. As shown in Tab. 9, FSFM achieves the best overall performance with a 0.75 masking ratio. Adopting lower masking ratios leads to trivial reconstruction and alignment tasks due to more available information. Conversely, using a higher masking ratio makes pretext tasks too challenging to learn sufficient facial representations for downstream face security tasks. Accordingly, we select a 75% masking ratio as the default setting.
Effect of Model Scaling Tab. 9 also shows that FSFM benefits from larger model sizes when pretrained on FF++_o. The transfer performance on downstream face security tasks improves as the model scales up. Due to limited computing resources, we were unable to pretrain larger models on more face images, but we aspire to explore this in future work.
Effect of Loss Weight To explore the impact of the reconstruction and distillation losses, we vary various loss weights, i.e. and . Results in Fig. 9 show that the configuration (, ) performs better on challenging cross-dataset deepfake detection.
F.2 Visualizations
Reconstruction To demonstrate the superiority of the facial representations pretrained with FSFM, we further follow MAE [45] to visualize reconstruction results, as shown in Fig. 10. We can see that FSFM demonstrates better reconstruction quality concerning intra-region consistency (preserving fine-grained textures within facial regions), inter-region coherency (maintaining spatial relationships across regions), and local-to-global correspondence (aligning local appearance with global facial looking).
CAM We provide additional CAM visualizations in Fig. 11, which are consistent with the observations in the main paper, further substantiating the effectiveness of our method.